# Vector Dark Matter with Higgs Portal in Type II Seesaw framework

### v1

@@ IV.1 Detection Aspects of DM
Dark matter detection can be classified broadly into three categories namely direct search, indirect search, and collider search. Direct search experiments measure the cross-section of elastic scattering between DM and nucleons inside the detector material. $Z^{\prime}$, in the present scenario, scatters off the nucleons via the mixing of $S$ with SM Higgs. The Feynman diagram of the elastic scattering of DM to nucleons is shown in Fig. 9. The approximate expression of the cross-section of this t-channel process is given by Baek:2012se
@@
where $\mu_{XN}=\frac{M_{N}M_{Z^{\prime}}}{M_{N}+M_{Z^{\prime}}}$ is reduced nucleon-DM mass and $f_{p}=\sum_{q=u,d,s}f^{p}_{q}+2/9(1-\sum_{q=u,d,s}f^{p}_{q})\simeq 0.468$ Belanger:2008sj . Here the effect of $H_{1}$ has been neglected in the formula of the DD cross-section. $H_{1}$, which is dominantly triplet, does have interaction with the nucleons or the DM suppressed by small factors involving mixing angles. As a result, the effective DM-nucleon scattering amplitude mediated by $H_{1}$ is proportional to $\sin\alpha_{1}\sin\alpha_{3}$. Therefore, in the small $\sin\alpha$ limit, the effect of $H_{1}$ can be neglected in the DM nucleon scattering cross-section.
@@
As given in the Eqn. 45, the direct search cross-section is sensitive to the parameters $M_{Z^{\prime}}$, $g_{x}$, $\sin\alpha_{2}$ and $m_{S}$. The product of $g_{x}$ and $\sin\alpha_{2}$ act as the effective coupling here. By imposing the collider bound on $\sin\alpha_{2}$, we can have an upper bound on the maximum allowed value of $g_{x}$ from DD limits. In Fig. 10, we show the bound on $g_{x}$ for a particular $\sin\alpha_{2}$ over a model parameter space satisfying all other constraints mentioned. In the left panel of Fig. 10, we show the relic density satisfied parameter space in $M_{Z^{\prime}}$ vs $g_{x}$ plane for a range of scalar masses allowed by vacuum stability and unitarity for $\sin\alpha_{2}=0.1$. It is observed that the maximum allowed value of $g_{x}$ is $\sim 0.3$ in this case from all constraints. The DD cross-section from the LZ experiment gives the most stringent bound and XENON constraints next to it. The other bounds include Higgs invisible decay width and the bound on $v_{s}$ though they are less stringent. In the right panel of Fig. 10, we show the parameter space for $\sin\alpha_{2}=0.01$ in the same plane. Here one can see that by lowering the value of $\sin\alpha_{2}$ by one order, the constraint on $g_{x}$ is significantly relaxed from DD.
@@
Figure 10: Parameter space with all mentioned constraints in $M_{Z^{\prime}}$ vs $g_{x}$ plane for $\sin\alpha_{2}=0.1$ in the left panel. The coloured region corresponds to exclusion from different constraints. In the right panel, same plot for $\sin\alpha_{2}=0.01$.
Indirect search provides a complementary way to look for DM in the universe. WIMP annihilation in the galactic center, nearby galaxies or the sun can produce observable fluxes of SM particles (photons, electrons, neutrinos etc.) which can be detected at various telescopes giving us information about DM distribution. Fermi-LAT, a satellite-based experimental facility, studies cosmic gamma-ray fluxes and provides the strongest bound on the DM annihilation to photon channel Goodenough:2009gk from the galactic centre up to few hundreds of GeV mass range. In this framework, the absence of tree-level coupling between the dark matter $(Z^{\prime})$ and photons still permits for the generation of photon line and diffuse signals of DM annihilation through these channels. Specifically, the processes $Z^{\prime}Z^{\prime}\to H_{2}\to\gamma\gamma$ at one loop and, if kinematically viable, $Z^{\prime}Z^{\prime}\to H_{2}\to W^{\pm}W^{\pm},H^{\pm}H^{\mp},H^{\pm\pm}H^{\mp\mp},f{\bar{f}}$ (with photons emitted from all external charged particles) are considered. However, the former channel is not only loop-induced but also subject to mixing angle suppression, while the latter channels are solely affected by mixing angle suppression. Consequently, due to the combined effects of loop and mixing angle suppression, both the line and diffuse gamma-ray fluxes resulting from dark matter annihilation in this case is significantly smaller than the limits indicated by Fermi-LAT data in the GeV mass ranges.
@@ V Possibility of Freeze-in
In the previous section, we explored a scenario where the DM was initially in thermal equilibrium with the SM particle bath, but at a later time, it froze out as the interaction rate dropped below the expansion rate of the universe ($\mathcal{H}$). Due to such interaction with the SM particles, it can be probed by different DD experiments as we have shown earlier. However, the null results of all the DD experiments motivate us to explore another possibility where the interaction of DM with SM is so feeble that it never reaches the thermal equilibrium in the early universe. Rather, it could be produced from the decay or annihilation of the bath particles. Due to such feeble interactions, it can evade the stringent DD constraints. In this model, $Z^{\prime}$ can gain its number density in the early universe from the decay of singlet scalar $S$ ($H_{2}$) or annihilation of $S$ ($H_{2}$). As $S$ can be thermally produced through their interaction with SM Higgs, if the gauge coupling of $U(1)_{X}$, $g_{x}$ is very small, typically in the order of ${}O(10^{-12})$, then $Z^{\prime}$ can never attain thermal equilibrium with the SM bath. With negligible initial abundance, DM gains its number density gradually from the decay of $S$. The decay of $S$ is dictated by the vertex factor corresponding to $S{Z^{\prime}}^{\mu}Z^{\prime}_{\mu}\sim g_{x}^{2}v_{S}$ whereas the annihilation of $S$ is controlled by the vertex factor of $S^{*}S{Z^{\prime}}^{\mu}Z^{\prime}_{\mu}\sim g_{x}^{2}$. As the annihilation of $S$ is suppressed by a $v_{s}$ factor with respect to the decay at the vertex factor level, the decay term gives a dominant contribution to DM production. The mixing of $S$ to other scalars being negligible $S$ can be identified with mass eigenstate $H_{2}$. To find out the DM number density evolution, we need to solve the Boltzmann equation for the number density of $Z^{\prime}$. The Boltzmann equation in terms of DM comoving number density $Y_{Z^{\prime}}$ is given by
@@
Now comoving number density of DM $Z^{\prime}$, $Y_{Z^{\prime}}$ can be analytically expressed as Hall:2009bx
@@
In the right panel of Fig. 11, we show a region in $M_{S}$ vs $M_{Z}^{\prime}$ plane where $Z^{\prime}$ can contribute in the relic density through the decay of scalar $S$. The colorbar corresponds to the percentage of relic density contribution of our DM candidate $Z^{\prime}$. The region under the allowed parameter space is overabundant where as the upper region is excluded by underabundance and kinamatics ($2M_{Z^{\prime}}\geq m_{S}$).
Before we conclude, let us very briefly comment on the search prospect of the VDM at the LHC. Unfortunately, the direct production of a $Z^{\prime}$ pair in pp collision is prohibited by conservation of $U(1)_{X}$ charge. $Z^{\prime}$ pair can arise from the decay of $H_{2}$ which can be produced in association with a $Z$ boson at the LHC. In the allowed range of model parameters ($m_{H_{2}},g_{x},\sin\alpha_{2}$) from the constraints mentioned, the cross-section is significantly small due to $\sin\alpha_{2}$ suppression. Another possible $Z^{\prime}$ signature can arise from the production of $H^{\pm\pm}$ in association with $H^{\mp}$ followed by the decays of $H^{\pm\pm}\rightarrow W^{\pm}W^{\pm}$ and $H^{\pm}\rightarrow W^{\pm}H_{2}$ and further followed by $H_{2}\rightarrow Z^{\prime}Z^{\prime}$. This final state would eventually result into two dilepton $+$ di-jet$+$missing energy signature. This cross-section is also very small due to the tiny $H^{+}\rightarrow W^{+}H_{2}$ suppressed branching ratio for allowed values of $m_{H_{2}}$ at the LHC. It seems that the detection of our VDM at the LHC is very challenging unless a more innovative method is divised.
@@ VI Conclusion
In this work, we study the phenomenology of a VDM in an $U(1)_{X}$ extension of the SM in the type II seesaw framework. The light neutrino masses have been generated via the type-II seesaw mechanism by introducing an additional Higgs triplet ($\Delta$). In the context of such a model, we investigate the possibility of a VDM, $Z^{\prime}$ which was stabilised by an exact dark conjugation symmetry. $Z^{\prime}$ becomes massive when a complex scalar singlet $S$ acquires VEV breaking the $U(1)_{X}$ symmetry spontaneously. We studied the possibility of both thermal and non-thermal production of $Z^{\prime}$ in the early universe. Thermalisation of $Z^{\prime}$ in the early universe can be achieved for a sizable $g_{x}$ through the interactions with SM particles via the mixing between SM Higgs and the newly added complex scalar. We have shown the implication of such WIMP DM in the light of the current limit of DD experiments. Due to the presence of the triplet scalar necessary for the type II seesaw mechanism, the vacuum stability and unitarity constraints on $\lambda$ parameters are conservative. Our model parameter space satisfies constraints from vacuum stability, perturbative unitarity, $\rho$ parameter, oblique parameters, lepton flavour violation from charged scalars and Higgs invisible decay. This framework can also accommodate the existing $2\sigma$ deviation observed in the $h\to Z\gamma$ over a wide range of parameter space. Here we show that after consideration of the recent LZ bound, huge parameter space corresponding to higher $g_{x}$ is excluded. However, such an upper limit on $g_{x}$ depends on the mixing angle between the SM Higgs and complex singlet scalar as shown in Fig. 10. At the end, we briefly discussed the possibility of non-thermal production of $Z^{\prime}$ from the decay of complex scalar $S.$
@@ Acknowledgement
Authors would like to thank Dilip Kumar Ghosh for various useful discussions during this work. ND would like to thank Pooja Bhattacharjee for useful discussions. ND is funded by CSIR, Government of India, under the NET SRF fellowship scheme with Award file No.09/080(1187)/2021-EMR-I. TJ acknowledges the support from the Science and Engineering Research Board (SERB), Government of India under grant reference no. PDF/2020/001053. The work of DN is supported by the National Research Foundation of Korea (NRF) Research Grant NRF-2019R1A2C3005009 (DN).
@@ A.2 $h\to\gamma\gamma$ and $h\to Z\gamma$ accessories
Here we can see that when $\sin\alpha_{2}$ tends to zero, we get back our pure type II seesaw couplings (see BhupalDev:2013xol ). In the limit, $\sin\alpha_{1}\to 0$, $\sin\alpha_{2}\to 0$ and $v_{t}<<v_{d}$, $\cos\beta^{\prime}\to 0$ and the relative couplings become one hence the SM Higgs coupling are restored. The scalar trilinear couplings $\tilde{g_{hH^{\pm}H^{\mp}}}$ and $\tilde{g_{hH^{\pm\pm}H^{\mp\mp}}}$ are given by
@@
The terms including $\lambda_{\Delta S}$ and $\lambda_{\Delta S}$ come from the extra complex singlet scalar sector and a factor of $\cos\alpha_{2}$ appears with the rest of the part due to its mixing with the SM doublet. In the limit $\sin\alpha_{2}\to 0$, these couplings reduce to Doublet Triplet Higgs model couplings as mentioned in Arhrib:2011vc . The necessary loop functions and couplings for $h\to Z\gamma$ are mentioned below. The loop factors for particles with different spins are given by
@@ A.3 Unitarty constraints
The unitary bounds of the model are extracted from the amplitude matrix ($2\to 2$ scattering amplitude ) where the basis eigen vector is made of all possible two particle states. Each eigen value of this matrix should be less than $8\pi$ to satisfy unitarity. The amplitude matrix can be decomposed into 9 charge neutral, 14 singly charged and 9 doubly charged two particle states. We have calculated the conditions following the prescription of Ref. Arhrib:2011uy which have been mentioned in the main text. However, the amplitude matrix corresponding to the charge-neutral scatterings does not give the closed form of eigenvalues and therefore we solve the matrix numerically. The charge neutral two particle states with identical particles are given by

### v2

@@ IV.1 Detection Aspects of DM
Dark matter detection can be classified broadly into three categories namely direct search, indirect search, and collider search. Direct search experiments measure the cross-section of elastic scattering between DM and nucleons inside the detector material. $Z^{\prime}$, in the present scenario, scatters off the nucleons via the mixing of $S$ with SM Higgs. The Feynman diagram of the elastic scattering of DM to nucleons is shown in Fig. 10. The approximate expression of the cross-section of this t-channel process is given by Baek et al. (2013)
@@
where $\mu_{XN}=\frac{M_{N}M_{Z^{\prime}}}{M_{N}+M_{Z^{\prime}}}$ is reduced nucleon-DM mass and $f_{p}=\sum_{q=u,d,s}f^{p}_{q}+2/9(1-\sum_{q=u,d,s}f^{p}_{q})\simeq 0.468$ Belanger et al. (2009). Here the effect of $H_{1}$ has been neglected in the formula of the DD cross-section. $H_{1}$, which is dominantly triplet, does have interaction with the nucleons or the DM suppressed by small factors involving mixing angles. As a result, the effective DM-nucleon scattering amplitude mediated by $H_{1}$ is proportional to $\sin\alpha_{1}\sin\alpha_{3}$. Therefore, in the small $\sin\alpha$ limit, the effect of $H_{1}$ can be neglected in the DM nucleon scattering cross-section.
@@
As given in the Eqn. 45, the direct search cross-section is sensitive to the parameters $M_{Z^{\prime}}$, $g_{x}$, $\sin\alpha_{2}$ and $m_{S}$. The product of $g_{x}$ and $\sin\alpha_{2}$ act as the effective coupling here. By imposing the collider bound on $\sin\alpha_{2}$, we can have an upper bound on the maximum allowed value of $g_{x}$ from DD limits. In Fig. 11, we show the bound on $g_{x}$ for a particular $\sin\alpha_{2}$ over a model parameter space satisfying all other constraints mentioned. In the left panel of Fig. 11, we show the relic density satisfied parameter space in $M_{Z^{\prime}}$ vs $g_{x}$ plane for a range of scalar masses allowed by vacuum stability and unitarity for $\sin\alpha_{2}=0.1$. It is observed that the maximum allowed value of $g_{x}$ is $\sim 0.3$ in this case from all constraints. The DD cross-section from the LZ experiment gives the most stringent bound and XENON constraints next to it. The other bounds include Higgs invisible decay width and the bound on $v_{s}$ from unitarity though their effect becomes important for smaller $\sin\alpha_{2}$. In the right panel of Fig. 11, we show the parameter space for $\sin\alpha_{2}=0.01$ in the same plane. Here one can see that by lowering the value of $\sin\alpha_{2}$ by one order, the constraint on $g_{x}$ is significantly relaxed from DD.
@@
Figure 11: Parameter space with all mentioned constraints in $M_{Z^{\prime}}$ vs $g_{x}$ plane for $\sin\alpha_{2}=0.1$ in the left panel. The coloured region corresponds to exclusion from different constraints. In the right panel, the same plot for $\sin\alpha_{2}=0.01$.
Indirect search provides a complementary way to look for DM in the universe. WIMP annihilation in the galactic center, nearby galaxies, or the sun can produce observable fluxes of SM particles (photons, electrons, neutrinos, etc.) which can be detected at various telescopes giving us information about DM distribution. Fermi-LAT, a satellite-based experimental facility, studies cosmic gamma-ray fluxes coming from the galactic centre. It provides the strongest bound on the rate of DM annihilation to photons for DM masses up to few hundreds of GeV Goodenough and Hooper (2009). In the present framework, the generation of monoenergetic and diffused photons from DM annihilation is driven by the processes $Z^{\prime}Z^{\prime}\to H_{2}\to\gamma\gamma$ at one loop level and via $Z^{\prime}Z^{\prime}\to H_{2}\to W^{\pm}W^{\pm},H^{\pm}H^{\mp},H^{\pm\pm}H^{\mp\mp},f{\bar{f}}$ (with photons emitted from all external charged particles) if kinematically possible. However, the process $Z^{\prime}Z^{\prime}\to H_{2}\to\gamma\gamma$ is not only loop-suppressed but also subject to mixing angle suppression. The other channels are solely affected by mixing angle suppression. Consequently, due to the combined effects of loop and mixing angle suppression, both the line and diffuse gamma-ray fluxes resulting from dark matter annihilation in our model are significantly smaller than the limits indicated by Fermi-LAT data in the GeV mass ranges. Other important constraints as mentioned in Elor et al. (2016) come from different experiments like CMB measurement by Planck, the Fermi detection of photons from the dwarf galaxies, the positron data from AMS-02, etc. However in our model, as DM annihilates to non-standard Higgses dominantly, so the direct production cross-section of SM particles is well below the observed bounds.
Before we conclude this section, let us very briefly comment on the search prospect of the VDM at the LHC. Unfortunately, the direct production of a $Z^{\prime}$ pair in pp collision is prohibited by conservation of $U(1)_{X}$ charge. $Z^{\prime}$ pair can arise from the decay of $H_{2}$ which can be produced in association with a $Z$ boson at the LHC. In the allowed range of model parameters ($m_{H_{2}},g_{x},\sin\alpha_{2}$) from the constraints mentioned, the cross-section is significantly small due to $\sin\alpha_{2}$ suppression. Another possible $Z^{\prime}$ signature can arise from the production of $H^{\pm\pm}$ in association with $H^{\mp}$ followed by the decays of $H^{\pm\pm}\rightarrow W^{\pm}W^{\pm}$ and $H^{\mp}\rightarrow W^{\mp}H_{2}$ and further followed by $H_{2}\rightarrow Z^{\prime}Z^{\prime}$. This final state would eventually result into two dilepton $+$ di-jet$+$missing energy signature. This cross-section is also very small due to the tiny $H^{\mp}\rightarrow W^{\mp}H_{2}$ suppressed branching ratio for allowed values of $m_{H_{2}}$ at the LHC. It seems that the detection of our VDM at the LHC is very challenging unless a more innovative method is devised.
@@ V Possibility of Freeze-in
In the previous section, we explored a scenario where the DM was initially in thermal equilibrium with the SM particle bath, but at a later time, it froze out as the interaction rate dropped below the expansion rate of the universe ($\mathcal{H}$). Due to such interaction with the SM particles, it can be probed by different DD experiments as we have shown earlier. However, the null results of all the DD experiments motivate us to explore another possibility where the interaction of DM with SM is so feeble that it never reaches the thermal equilibrium in the early universe. Rather, it could be produced from the decay or annihilation of the bath particles. Small value of coupling necessary for such feeble interaction can be naturally generated in a number of ways as shown in Ref. Babu et al. (2023) in a flavor model context. Due to such feeble interactions, DM can evade the stringent DD constraints and collider constraints. In this model, $Z^{\prime}$ can gain its number density in the early universe from the decay of singlet scalar $S$ ($H_{2}$) or annihilation of $S$ ($H_{2}$). As $S$ can be thermally produced through its interaction with SM Higgs, if the gauge coupling of $U(1)_{X}$, $g_{x}$ is very small, typically in the order of ${}O(10^{-12})$, then $Z^{\prime}$ can never attain thermal equilibrium with the SM bath. With negligible initial abundance, DM gains its number density gradually from the decay of $S$. The decay of $S$ is dictated by the vertex factor corresponding to $S{Z^{\prime}}^{\mu}Z^{\prime}_{\mu}\sim g_{x}^{2}v_{S}$ whereas the annihilation of $S$ is controlled by the vertex factor of $S^{*}S{Z^{\prime}}^{\mu}Z^{\prime}_{\mu}\sim g_{x}^{2}$. As the annihilation of $S$ is suppressed by a $v_{s}$ factor with respect to the decay at the vertex factor level, the decay term gives a dominant contribution to DM production. The mixing of $S$ to other scalars being negligible $S$ can be identified with mass eigenstate $H_{2}$. To find out the DM number density evolution, we need to solve the Boltzmann equation for the number density of $Z^{\prime}$. The Boltzmann equation in terms of DM comoving number density $Y_{Z^{\prime}}$ is given by
@@
Now comoving number density of DM $Z^{\prime}$, $Y_{Z^{\prime}}$ can be analytically expressed as Hall et al. (2010)
@@
In the right panel of Fig. 12, we show a region in $M_{S}$ vs $M_{Z}^{\prime}$ plane where $Z^{\prime}$ can contribute to the relic density through the decay of scalar $S$ for $g_{x}=10^{-12}$. The color bar corresponds to the percentage of relic density contribution of our DM candidate $Z^{\prime}$. The region under the allowed parameter space is overabundant whereas the upper region is excluded by kinematics ($2M_{Z^{\prime}}\geq m_{S}$).
@@ VI Conclusion
To summarise, we have studied the phenomenology of a VDM in a gauged $U(1)_{X}$ extension of the SM set in a type II seesaw framework. The light neutrino masses are generated via the type-II seesaw mechanism by the introduction of an additional Higgs triplet ($\Delta$). In the context of such a model, we investigate the possibility of a VDM, $Z^{\prime}$ which is stabilised by an exact dark conjugation symmetry. $Z^{\prime}$ becomes massive when a complex scalar singlet $S$ acquires VEV breaking the $U(1)_{X}$ symmetry spontaneously. We studied the possibility of both thermal and non-thermal production of $Z^{\prime}$ in the early universe. Thermalisation of $Z^{\prime}$ in the early universe can be achieved for a sizable range of values of $g_{x}$ ($0-1$) along with a comparatively smaller value of $\lambda_{s\phi}$ ($\mathcal{O}(10^{-4})$), the mixing between SM Higgs and the additional singlet scalar ($S$). This mixing acts as a portal between the dark sector and the visible sector. To begin with, we have estimated the density of $Z^{\prime}$ in the present epoch by implementing our model in micromega. We have then shown a comparison between the relic density satisfied parameter space of our scenario and that of the $U(1)_{X}$ extended minimal VDM model. We found out that our scenario allows lower $g_{x}$ values for larger mass of DM due to presence of newer annihilation channels with respect to the minimal model. The resulting relic density has been compared with available data reported by the PLANCK satellite. We have shown the implications of such a WIMP DM in light of the current limits of DD experiments like XENON and LZ. Due to the presence of the triplet scalar, necessary for neutrino mass generation, the vacuum stability and unitarity constraints on $\lambda$ parameters are restrictive. Our model parameter space satisfies constraints from vacuum stability, perturbative unitarity, $\rho$ parameter, oblique parameters, lepton flavour violation from charged scalars, and Higgs invisible decay. This framework can also accommodate the existing $2\sigma$ deviation observed in the $h\to Z\gamma$ decay rate over a wide range of parameter space. We have shown that after considering all the aforementioned constraints, the upper limit of $g_{x}$ is nearly $0.3$ for DM masses in the range $(100-1000)$ GeV when $\sin\alpha_{2}=0.1$. However, such an upper limit on $g_{x}$ crucially depends on $\sin\alpha_{2}$ and becomes weaker for lower values of $\sin\alpha_{2}$ as evident from Fig. 11. In the end, we briefly discussed the possibility of non-thermal production of $Z^{\prime}$ from the decay of the complex scalar $S$. For the values of $g_{x}$ as small as $10^{-12}$, the allowed mass range of DM varies between $(100-300)$ GeV when $Z^{\prime}$ contributes $100\%$ to the relic density.
@@ Acknowledgement
The authors would like to thank Dilip Kumar Ghosh for various useful discussions during this work. ND would like to thank Pooja Bhattacharjee for useful discussions. ND is funded by CSIR, Government of India, under the NET SRF fellowship scheme with Award file No.09/080(1187)/2021-EMR-I. TJ acknowledges the support from the Science and Engineering Research Board (SERB), Government of India under grant reference no. PDF/2020/001053. The work of DN is supported by the National Research Foundation of Korea (NRF) Research Grant NRF-2019R1A2C3005009 (DN).
@@ A.2 $h\to\gamma\gamma$ and $h\to Z\gamma$ accessories
Here we can see that when $\sin\alpha_{2}$ tends to zero, we get back our pure type II seesaw couplings (see Bhupal Dev et al. (2013)). In the limit, $\sin\alpha_{1}\to 0$, $\sin\alpha_{2}\to 0$ and $v_{t}<<v_{d}$, $\cos\beta^{\prime}\to 0$ and the relative couplings become one hence the SM Higgs coupling are restored. The scalar trilinear couplings $\tilde{g_{hH^{\pm}H^{\mp}}}$ and $\tilde{g_{hH^{\pm\pm}H^{\mp\mp}}}$ are given by
@@
The terms including $\lambda_{\Delta S}$ and $\lambda_{\Delta S}$ come from the extra complex singlet scalar sector and a factor of $\cos\alpha_{2}$ appears with the rest of the part due to its mixing with the SM doublet. In the limit $\sin\alpha_{2}\to 0$, these couplings reduce to Doublet Triplet Higgs model couplings as mentioned in Arhrib et al. (2012). The necessary loop functions and couplings for $h\to Z\gamma$ are mentioned below. The loop factors for particles with different spins are given by
@@ A.3 Unitarty constraints
The unitary bounds of the model are extracted from the amplitude matrix ($2\to 2$ scattering amplitude ) where the basis eigen vector is made of all possible two particle states. Each eigen value of this matrix should be less than $8\pi$ to satisfy unitarity. The amplitude matrix can be decomposed into 9 charge neutral, 14 singly charged and 9 doubly charged two particle states. We have calculated the conditions following the prescription of Ref. Arhrib et al. (2011) which have been mentioned in the main text. However, the amplitude matrix corresponding to the charge-neutral scatterings does not give the closed form of eigenvalues and therefore we solve the matrix numerically. The charge neutral two particle states with identical particles are given by